OpenStudy (anonymous):

Let f(x)=6^x. Which of the following is NOT true about the graph of y=f(x)?

OpenStudy (anonymous):

(a) The graph of y=f(x) lies above the x-axis (b) The graph of y=f(x) is always increasing (c) The point (1,6) is on the graph of y=f(x) (d) The graph of y=f(x) has an y-intercept at the point (0,1) (e) The graph of y=f(x) has an x-intercept at the point (1,0)

OpenStudy (anonymous):

a) is true, it is always positive b) is true it is always increasing c) is true because 6^1=6 d) is true because 6^0=1 i guess only wrong one is e, since 6^1 is not 0

OpenStudy (anonymous):

e